A bottle can hold 900 mL of liquid when it is 2/3 full. How many mL of liquid can the bottle hold if it is full?

Answer by stanbon(75887)

A bottle can hold 900 mL of liquid when it is 2/3 full. How many mL of liquid can the bottle hold if it is full?
-------
Let the total amount it can hold be "x"::
Equation:
(2/3)x = 900
x = (3/2)900
----
x = 3(450)
x = 1350 mL
--------------
Cheers,
Stan H.
-----------